### 24 Mar 2026 v1

# Classify all representation which contains a Steinberg in its hyperspecial subgroup

Abstract
This paper addresses Question 1 posed by Dipendra Prasad in his recent problem list: classify all irreducible smooth representations $\pi$ of an unramified reductive $p$-adic group $G$ (over a characteristic zero local field) such that the space of $K_{+}$-fixed vectors $V^{K_{+}}$ (viewed as a representation of the finite reductive group $K/K_{+}$) contains the Steinberg representation. We prove that any such representation $\pi$ must be Iwahori-spherical, hence a subquotient of some unramified principal series $I(\chi)=\operatorname{Ind}_{B}^{G}\chi$. By a detailed analysis of the action of the Iwahori–Hecke algebra on the Iwahori-fixed space $V^{I}$, we show that in each principal series $I(\chi)$ there exists exactly one irreducible subquotient containing Steinberg, and we give an explicit construction of this subquotient. Hence this can give us a bijection between W-orbits of unramified characters and irreducible representations containing Steinberg in its hyperspecial subgroup. This classification is richer than Prasad’s expectation: besides twists of the Steinberg representation and generic unramified representations, every principal series contributes a unique subquotient with the required property. The proof relies on establishing an isomorphism between the finite Iwahori–Hecke algebra of the finite Lie group $K/K_{+}$ and the finite Hecke algebra $H(K,I)$ of the $p$-adic group (compactly supported $I$-bi-invariant functions on $K$), and comparing their actions on $V^{I}$.
Keywords: $p$-adic groups, unramified reductive groups, Steinberg representation, Iwahori subgroup, Hecke algebras, unramified principal series, Moy–Prasad filtration.
1 Introduction
Let $F$ be a non-archimedean local field of characteristic zero, $\mathcal{O}_{F}$ its ring of integers, $\mathbb{F}_{q}$ its residue field, and let $G$ be a connected reductive group defined over $F$ which is unramified, i.e., quasi-split and split over an unramified extension of $F$. Denote by $K=G(\mathcal{O}_{F})$ a hyperspecial maximal compact subgroup of $G(F)$ and by $K_{+}$ its pro-unipotent radical; then $K/K_{+}\simeq G(\mathbb{F}_{q})$ is a finite reductive group. Through inflation, representations of $G(\mathbb{F}_{q})$ may be viewed as representations of $K$.
In a recent discussion meeting at IISER Bhopal (July 2025), Dipendra Prasad proposed a list of accessible open problems in representation theory [10]. The first problem reads:
Classify all irreducible representations of $G(F)$ containing the Steinberg representation of $G(\mathbb{F}_{q})$ as a $K$-subrepresentation (i.e., such that $V^{K_{+}}$ contains the Steinberg representation of $K/K_{+}$ as a constituent).
Prasad suggested that such representations should fall into two families: (1) twists of the Steinberg representation of $G(F)$; (2) generic, unramified representations (i.e., irreducible spherical principal series). However, our investigation reveals a more nuanced picture: every unramified principal series $I(\chi)=\operatorname{Ind}_{B}^{G}\chi$ (where $B$ is a Borel subgroup and $\chi$ an unramified character of a maximal torus) contains exactly one irreducible subquotient that contains the Steinberg representation. In particular, the class of representations satisfying Prasad’s condition is strictly larger than his expectation; each principal series contributes a unique such subquotient, hence this can give us a bijection between W-orbits of unramified characters and irreducible representations containing Steinberg in its hyperspecial subgroup.
The main steps of our argument are as follows. Let $(\pi,V)$ be an irreducible smooth representation of $G(F)$ with $V^{K_{+}}$ containing the Steinberg representation. By the depth-zero theory of Moy–Prasad [8, 9], $\pi$ must be of depth zero. Using Lemma 3.2 and the Borel–Casselman theorem [1], we deduce that $\pi$ is Iwahori-spherical, i.e., $V^{I}\neq 0$ for an Iwahori subgroup $I\subset K$. Hence $\pi$ is a subquotient of some unramified principal series $I(\chi)$. The space $V^{K_{+}}$ is a representation of the finite reductive group $K/K_{+}$; by Lemma 3.1 all its irreducible constituents lie in the principal Harish-Chandra series, i.e., are subrepresentations of $\operatorname{Ind}_{I/K_{+}}^{K/K_{+}}\mathbf{1}$. Taking $I/K_{+}$-fixed points yields $V^{I}$, which carries two natural actions: one from the Hecke algebra $H_{\mathrm{fin}}=\operatorname{End}_{K/K_{+}}(\operatorname{Ind}_{I/K_{+}}^{K/K_{+}}\mathbf{1})$ (coming from finite group theory), and one from the finite Hecke algebra $H(K,I)$ of the $p$-adic group (compactly supported $I$-bi-invariant functions on $K$). In Section 6 we prove that these two actions coincide on $V^{I}$ (Lemma 6.2), allowing us to transfer known results about the Steinberg representation from finite groups to the $p$-adic setting.
The multiplicity of the Steinberg representation in $V^{K_{+}}$ equals the dimension of the space
$W=\{v\in V^{I}\mid T_{\alpha}v=-v\text{ for\ every \ simple \ root \ }\alpha\},$
where $T_{\alpha}$ are the standard generators of the Hecke algebra. Using Casselman’s basis $\{\phi_{w,\chi}\}$ of $I(\chi)^{I}$ and Lansky’s formulas for the action of affine reflections [7] (adapted to unnormalized induction), we solve this eigenvalue problem in $I(\chi)^{I}$. Lemma 7.2 shows that the solution space is one‑dimensional; a non‑zero vector is explicitly given by
$v_{0}=\sum_{w\in W}(-1)^{\ell(w)}[IwI:I]^{-1}\,\phi_{w,\chi}.$
Consequently $\dim W=1$, so the Steinberg representation occurs with multiplicity one in $V^{K_{+}}$. Moreover, letting $V$ be the subrepresentation of $I(\chi)$ generated by $v_{0}$ and $V_{0}$ the maximal subrepresentation of $V$ not containing $v_{0}$, the quotient $V/V_{0}$ is the unique irreducible subquotient of $I(\chi)$ that contains Steinberg. This completes the classification.
The paper is organized as follows. Section 2 recalls the Moy–Prasad filtration and unrefined minimal $K$-types. Section 3 analyses the cuspidal support of the irreducible components of $V^{K_{+}}$ and proves that a representation containing Steinberg must be Iwahori-spherical. Section 4 collects facts about unramified principal series: Casselman’s basis and Lansky’s action of affine reflections. Section 5 reviews the representation theory of finite Lie groups, focusing on the principal series and the associated Iwahori–Hecke algebra. Section 6 establishes the isomorphism between the finite Hecke algebra and the $p$-adic finite Hecke algebra $H(K,I)$, and proves that their actions on $V^{I}$ coincide. Section 7 contains the core lemma and the proof of the main theorem.
2 Unrefined minimal $K$-types and the Moy–Prasad filtration
In this section we recall the notion of Moy–Prasad filtrations and unrefined minimal $K$-types for a reductive $p$-adic group, following [8, 9]. Let $k$ be a non-archimedean local field of characteristic zero, and let $G$ be a connected reductive group defined over $k$. Let $\mathfrak{g}$ be the Lie algebra of $G$ and $\mathfrak{g}^{*}$ its dual.
2.1 Moy–Prasad filtrations
For a point $x$ in the Bruhat–Tits building of $G/k$, let $\mathcal{P}_{x}=\mathcal{P}_{x}\cap G(k)$ be the corresponding parahoric subgroup of $G(k)$. Moy and Prasad associate to $x$ a decreasing filtration $\{\mathcal{P}_{x,r}\}_{r\geq 0}$ of $\mathcal{P}_{x}$ by open compact subgroups, defined as follows: for a non-negative real number $r$, let $\mathcal{P}_{x,r}$ be the subgroup generated by certain congruence subgroups of the maximal torus and by the root subgroups $U_{\psi}$ for all affine roots $\psi$ satisfying $\psi(x)\geq r$; see [8, 2.6] for the precise definition. One sets $\mathcal{P}_{x,0}=\mathcal{P}_{x}$ and $\mathcal{P}_{x,r+}=\bigcup_{s>r}\mathcal{P}_{x,s}$. For $r>0$ the quotients $\mathcal{P}_{x,r}/\mathcal{P}_{x,r+}$ are abelian.
There is an associated filtration of the Lie algebra $\mathfrak{g}$. Let $\mathfrak{g}_{x,r}$ be the $\mathfrak{o}$-lattice in $\mathfrak{g}$ obtained from the smooth $\mathfrak{o}$-group scheme corresponding to $\mathcal{P}_{x}$; its definition is given in [9, 3.2]. Similarly, one defines a filtration $\{\mathfrak{g}_{x,r}^{*}\}_{r\in\mathbb{R}}$ of the dual $\mathfrak{g}^{*}$ by
$\mathfrak{g}_{x,r}^{*}=\{X\in\mathfrak{g}^{*}\mid X(Y)\in\mathfrak{o}\text{ for all }Y\in\mathfrak{g}_{x,-r+}\},$
see [8, 3.5]. The pairing
$\mathfrak{g}_{x,-r}^{*}/\mathfrak{g}_{x,(-r)+}^{*}\times\mathfrak{g}_{x,r}/\mathfrak{g}_{x,r+}\to\mathfrak{f}$
is non-degenerate and $\mathbf{M}_{x}(\mathfrak{f})$-equivariant, where $\mathbf{M}_{x}$ is the reductive quotient of the special fibre of the integral model attached to $x$; cf. [9, 3.7].
For a character $\chi$ of $\mathcal{P}_{x,r}$ trivial on $\mathcal{P}_{x,r+}$ (with $r>0$), one can realize $\chi$ as a character of the abelian group $\mathcal{P}_{x,r}/\mathcal{P}_{x,r+}$. Via the isomorphism $\mathcal{P}_{x,r}/\mathcal{P}_{x,r+}\cong\mathfrak{g}_{x,r}/\mathfrak{g}_{x,r+}$ of [9, 3.8] and the above duality, such a character corresponds to a coset
$X+\mathfrak{g}_{x,-r+}^{*}\in\mathfrak{g}_{x,-r}^{*}/\mathfrak{g}_{x,-r+}^{*}.$
Definition 2.1.
[8] Let $r>0$. A character $\chi$ of $\mathcal{P}_{x,r}$ trivial on $\mathcal{P}_{x,r+}$ is called non-degenerate if the corresponding coset $X+\mathfrak{g}_{x,-r+}^{*}$ contains no nilpotent elements.
2.2 Unrefined minimal $K$-types
Definition 2.2 ([8], 5.1).
An unrefined minimal $K$-type is a pair $(\mathcal{P}_{x,r},\chi)$ where $x$ is a point of the Bruhat–Tits building of $G/k$, $r$ is a non-negative real number, and $\chi$ is a representation of $\mathcal{P}_{x,r}$ trivial on $\mathcal{P}_{x,r+}$ such that:
  • (i)
    if $r=0$, then $\chi$ is a cuspidal representation of $\mathcal{P}_{x,0}/\mathcal{P}_{x,0+}$, inflated to $\mathcal{P}_{x}$;
  • (ii)
    if $r>0$, then $\mathcal{P}_{x,r}\neq\mathcal{P}_{x,r+}$ and $\chi$ is a non-degenerate character of the abelian group $\mathcal{P}_{x,r}/\mathcal{P}_{x,r+}$.
The number $r$ is called the depth of the minimal $K$-type.
Definition 2.3 ([9]).
Two unrefined minimal $K$-types $(\mathcal{P}_{x,r},\chi)$ and $(\mathcal{P}_{y,s},\xi)$ are said to be associates if they have the same depth $r=s$ and the following conditions hold:
  • (i)
    Zero depth case: $r=s=0$. Then there exists an element $g\in G(k)$ such that $\mathcal{P}_{x}\cap g\mathcal{P}_{y}g^{-1}$ surjects onto both $\mathbf{M}_{x}(\mathfrak{f})$ and $\mathbf{M}_{gy}(\mathfrak{f})$, and the representation $\chi$ is isomorphic to $\operatorname{Ad}(g)\xi$.
  • (ii)
    Positive depth case: $r=s>0$. Let $X+\mathfrak{g}_{x,-r+}^{*}$ be the coset representing $\chi$ and $Y+\mathfrak{g}_{y,-r+}^{*}$ the coset representing $\xi$. Then the $G(k)$-orbit of $X+\mathfrak{g}_{x,-r+}^{*}$ intersects $Y+\mathfrak{g}_{y,-r+}^{*}$.
2.3 The main theorem on existence and depth
The following fundamental result establishes the existence of an unrefined minimal $K$-type in any irreducible admissible representation, the uniqueness of its depth, and the associativity of all minimal $K$-types occurring in the same representation.
Theorem 2.4 ([8]).
Assume that $k$ is of characteristic zero. Let $(\pi,V_{\pi})$ be an irreducible admissible complex representation of $G(k)$. Then there exists a non-negative rational number $\rho(\pi)$ with the following properties:
  1. (1)
    For some point $x$ in the Bruhat–Tits building of $G/k$, the space $V_{\pi}^{\mathcal{P}_{x,\rho(\pi)+}}$ is non-zero, and $\rho(\pi)$ is the smallest number with this property.
  2. (2)
    For any point $y$ in the building, if $W=V_{\pi}^{\mathcal{P}_{y,\rho(\pi)+}}\neq\{0\}$, then:
    • (i)
      if $\rho(\pi)=0$, any irreducible $\mathcal{P}_{y,0}$-submodule of $W$ contains a minimal $K$-type of depth zero of a parahoric subgroup $\mathcal{Q}\subset\mathcal{P}_{y}$;
    • (ii)
      if $\rho(\pi)>0$, any irreducible $\mathcal{P}_{y,\rho(\pi)}$-submodule of $W$ is a minimal $K$-type.
Moreover, any two minimal $K$-types contained in $\pi$ are associates of each other.
The number $\rho(\pi)$ is called the depth of the representation $\pi$. In particular, every depth-zero irreducible representation contains an unrefined minimal $K$-type.
3 The cuspidal support of the irreducible components of $K/K_{+}$
Lemma 3.1.
Let $G$ be an unramified reductive $p$-adic group, let $K$ be a hyperspecial subgroup of $G$, and let $K_{+}$ be its corresponding pro-unipotent radical. Let $(\pi,V)$ be an irreducible smooth depth-zero representation of $G$. Then we may regard $V^{K_{+}}$ naturally as a representation of $K$ and hence also as a representation of the finite reductive group $K/K_{+}$. Under these assumptions, then for an arbitrary irreducible $K/K_{+}$-subrepresentation $W$ of $V^{K_{+}}$, there exists a facet $c$ and a cuspidal representation $\rho$ of the finite Lie group $G_{c}/G_{c+}$, such that the vertex associated to $K$ is contained in the closure of $c$ and W contains $\rho$. (We can also prove that $c,\rho]$ is a unrefined minimal K-type of V ). Moreover, every irreducible component of the $K/K_{+}$ representation is a subrepresentation of a Harish-Chandra series of the finite Lie group $K/K_{+}$.
Proof.
Since $V$ has depth zero by hypothesis, the space $V^{K_{+}}$ is non-zero.
Let $W$ be an irreducible $K/K_{+}$-subrepresentation of $V^{K_{+}}$.
By Theorem 2.4(2) , $W$ contains a depth-zero minimal $K$-type associated to a parahoric subgroup $G_{c}\subseteq K$ (where $c$ is a facet in the Bruhat–Tits building $\mathcal{B}(G)$). Denote this $G_{c}/G_{c+}$-cuspidal representation by $\rho$. The subgroup $G_{c}/K_{+}$ is a parabolic subgroup of $K/K_{+}$ with Levi factor $G_{c}/G_{c+}$. For simplicity, write this Levi decomposition as $Q=LU$. Where $L$ is the Levi factor and $U$ is its unipotent radical.
There exists a $Q$-embedding of $\rho$ into $W$. Since $\rho$ is $U$-invariant, the image of this embedding lies in $W^{U}$. Note that
$W^{U}\subset(V^{K_{+}})^{G_{c+}/K_{+}}=V^{G_{c+}}.$
Because $\rho$ is cuspidal, the pair $\rho,c]$ is an unrefined minimal $K$-type of $V$ by the definition of the unrefined minimal K-type .
Moreover, we can see that every irreducible component of $V^{K_{+}}$ is a subrepresentation of the parabolic induction from a cuspidal representation of a Levi (this is Harish-Chandra induction since we can find the parabolic and we need not use Deligne-Lusztig induction). ∎
To simplify terminology, if the representation $(\pi,V)$ of $G$ satisfies that the $K/K_{+}$-representation $V^{K_{+}}$ contains a Steinberg representation, we say that $V$ contains Steinberg.
Lemma 3.2.
Let $G$ be as in Lemma 3.1, let $I$ be its Iwahori subgroup, and let $(\pi,V)$ be an irreducible representation of $G$. If $V$ contains Steinberg, then $V$ must have a non-zero $I$-fixed vector.
Proof.
If $V$ contains Steinberg, then there exists an irreducible constituent of $V^{K_{+}}$ isomorphic to the Steinberg representation. The cuspidal support of the Steinberg representation is $(T,\mathbf{1})$, where $T$ is a torus and $\mathbf{1}$ denotes the trivial representation.
By Lemma 3.1, there exists a chamber $c$ whose closure contains the vertex associated to $K$ (Assume the Iwahori subgroup associated to $c$ is $I$ ), such that Steinberg contains this trivial representation, and the unrefined minimal $K$-type of $V$ is $(c,\mathbf{1})$. This means that the $I/I_{+}$-representation $V^{I_{+}}$ must contain a trivial representation; hence $V^{I}\neq 0$. ∎
A classical result of Borel and Casselman relates the existence of Iwahori fixed vectors to principal series representations.
Theorem 3.3 (Borel–Casselman, [1]).
Let $G$ be a connected reductive group over a non-archimedean local field $k$, and let $I\subset G(k)$ be an Iwahori subgroup. For an irreducible smooth representation $\pi$ of $G(k)$, the following are equivalent:
  1. 1.
    $\pi^{I}\neq 0$;
  2. 2.
    $\pi$ is isomorphic to a subquotient of an unramified principal series representation of $G(k)$.
Combining Theorem 3.3 and Lemma 3.2, we obtain the following immediate corollary.
Corollary 3.4.
Assume that $G$ satisfies the hypotheses of Lemma 3.1. Let $(\pi,V)$ be an irreducible representation of $G$. If $V$ contains Steinberg, then $\pi$ must be a subquotient of an unramified principal series.
Lemma 3.5.
Suppose $G$ is an unramified reductive $p$-adic group, and $(\pi,V)$ is an Iwahori-spherical irreducible representation of $G$. And fix an Iwahori subgroup $I$ which is contained in K. Then every irreducible $K/K_{+}$-subrepresentation $W$ of $V^{K_{+}}$ lies in the principal Harish-Chandra series $\operatorname{Ind}_{I/K_{+}}^{K/K_{+}}(1)$ (i.e., parabolic induction from the trivial representation of a quasi-split torus).
Proof.
For a fixed Iwahori subgroup $I$ of $G$ , By [6, Theorem 8.4.10(2)], we know that $I/I_{+}$ is the minimal Levi of $K/K_{+}$. Since every finite Lie group is quasi-split, $I/I_{+}$ is a torus, and $I/K_{+}$ is the Borel containing this torus.
Since $V$ is Iwahori-spherical, $V^{I}\neq 0$; hence the unrefined minimal $K$-type of $V$ is (the chamber corresponding to $I$, trivial representation). Since two unrefined minimal K type is associated, every unrefined minimal K type of V has the form (chamber, trivial representation).
By the proof of Lemma 3.1, we know that for every irreducible component $W$ of $V^{K_{+}}$ , there exists a chamber c (assume the Iwahori subgroup associated to c is I’), such that W is a subrepresentation of $\operatorname{Ind}_{I^{\prime}/K_{+}}^{K/K_{+}}(1)$. It is well known that every Borel subgroups of a finite Lie group are rational-conjugated. Hence $\operatorname{Ind}_{I^{\prime}/K_{+}}^{K/K_{+}}(1)$ is isomorphic to $\operatorname{Ind}_{I/K_{+}}^{K/K_{+}}(1)$. Hence the lemma holds. ∎
4 Some facts about the unramified principal series
4.1 Iwahori fixed subspace of an unramified principal series
The material in this subsection is derived from the seminal work of Casselman [2]. Let $G$ be an unramified, quasi-split, connected reductive group defined over a local field $k$. Let $B$ be a Borel subgroup of $G$ with Levi subgroup $T$. We denote by $K$ a hyperspecial subgroup of $G$ and by $I$ the Iwahori subgroup contained in $K$. Let $\chi$ be an unramified character of $T$. We define $I(\chi)$ as the unnormalized induced representation $\operatorname{Ind}_{B}^{G}\chi$.
In this subsection, we exhibit a canonical basis for the Iwahori fixed subspace $I(\chi)^{I}$. Proposition 2.1 of Casselman’s paper [2] states that the dimension of $I(\chi)^{I}$ equals the order of the relative Weyl group of $G$. We can construct these functions more concretely. To proceed, we first recall some notation.
Let $C_{c}^{\infty}(G)$ denote the space of compactly supported, locally constant, complex-valued functions on $G$. We define a projection from $C_{c}^{\infty}(G)$ onto $I(\chi)$. For a function $f\in C_{c}^{\infty}(G)$, we define its projection onto $I(\chi)$ by the formula
$\mathcal{P}_{\chi}(f)(g)=\int_{B}\chi^{-1}(b)f(bg)\,db,$
where $db$ is a left-invariant Haar measure on the Borel subgroup $B$ normalized so that $\operatorname{vol}(K\cap B)=1$.
For any subset $X\subset G$, let $\operatorname{ch}_{X}$ denote its characteristic function. Let $W$ be the relative Weyl group of $G$. In [2, Proposition 2.1], Casselman showed that the functions $\mathcal{P}_{\chi}(\operatorname{ch}_{BwB})$, as $w$ runs over all elements of $W$, form a basis for the vector space $I(\chi)^{I}$. We denote $\mathcal{P}_{\chi}(\operatorname{ch}_{BwB})$ by $\phi_{w,\chi}$. We formally state this result as follows.
Theorem 4.1 ([2]).
The functions $\phi_{w,\chi}$ for $w\in W$ form a basis of $I(\chi)^{I}$.
To provide a more explicit description, we can write down the values of these functions concretely. For any $b\in B$, $w\in W$, and $i\in I$, we have
$\phi_{w,\chi}(bwi)=\chi(b).$
Furthermore, the support of $\phi_{w,\chi}$ is contained in the double coset $BwI$.
It is well known that the $K$-fixed subspace $I(\chi)^{K}$ is one-dimensional. It is not hard to see that the sum $\sum_{w\in W}\phi_{w,\chi}$ is a non-zero vector in $I(\chi)^{K}$.
4.2 Effect of affine reflections on the Iwahori-fixed vectors
In this section we recall the results of Lansky [7] on the action of the affine Weyl group on the space of Iwahori-fixed vectors of an unramified principal series representation. We adapt his notation to match our conventions: let $I$ denote an Iwahori subgroup (denoted by $B$ in [7]), and let $B$ denote a minimal parabolic subgroup (denoted by $P$ in [7]). The opposite parabolic is denoted $B^{-}$. We use unnormalized induction, i.e., the induced representation $\operatorname{Ind}_{B}^{G}\chi$ without the modular factor $\delta^{1/2}$; consequently, the formulas in Corollary 3.2 of [7] simplify by omitting all occurrences of $\delta^{1/2}$. The maximal split torus is denoted by $S$ (instead of $A$ in [7]).
We briefly recall the notation used in [7] which will be needed in the sequel. Let $G$ be a quasi-split connected reductive $p$-adic group, and fix a maximal split torus $S$ contained in a minimal parabolic $B$. And assume that the root system of G relative to S is $\Phi^{\prime}$ and suppose that $\Phi$ is the sub root system vanishing on the hyperspecial point $x_{0}$. $\Phi$ is reduced and has a bijection with $^{nd}\Phi^{\prime}$. Let $M$ be the Levi component of $B$ (the centralizer of $S$), and let $N$ be the unipotent radical of $B$. Write $B=MN$. Let $B^{-}=MN^{-}$ be the opposite parabolic. From now on, we normalize the Haar measure on $G$ such that $I$ has volume 1.
Let $W=N_{G}(S)/M$ be the relative Weyl group. Let $\mathcal{N}$ be the normalizer of $S$ in $G$, and let $M_{0}$ be the maximal compact subgroup of $M$. Define the Iwahori–Weyl group $\widetilde{W}=\mathcal{N}/M_{0}$. There is a natural surjection $\nu:\widetilde{W}\to W$. The affine Weyl group $W_{\mathrm{aff}}$ is a normal subgroup of $\widetilde{W}$ generated by affine reflections.
Let $\Phi$ be the reduced root system obtained from the affine roots vanishing at a fixed special point $x_{0}$; it is identified with the set of non-divisible relative roots. Let $\Phi^{+}$ be the positive roots determined by $B$, and let $\Delta$ be the corresponding set of simple roots. For each irreducible component $\Phi_{i}$ of $\Phi$, let $\widetilde{\alpha}_{i}$ denote its highest root, and let $\widetilde{\Delta}$ be the set of these highest roots.
The affine root system $\Phi_{\mathrm{aff}}$ consists of the affine functions $\alpha+k$ ($\alpha\in\Phi$, $k\in\mathbb{Z}$) and $\widetilde{\alpha}-1$ ($\widetilde{\alpha}\in\widetilde{\Delta}$). The set of simple affine roots is
$\Delta_{\mathrm{aff}}=\{\alpha\mid\alpha\in\Delta\}\cup\{\widetilde{\alpha}-1\mid\widetilde{\alpha}\in\widetilde{\Delta}\}$.
For each $\alpha\in\Delta_{\mathrm{aff}}$, let $w_{\alpha}$ be the corresponding affine reflection in $W_{\mathrm{aff}}$. The set $S_{\mathrm{aff}}=\{w_{\alpha}\mid\alpha\in\Delta_{\mathrm{aff}}\}$ generates $W_{\mathrm{aff}}$ as a Coxeter group.
For $\alpha\in\Phi$, define the translation $a_{\alpha}=w_{\alpha}w_{\alpha-1}\in W_{\mathrm{aff}}$. One has $a_{-\alpha}=a_{\alpha}^{-1}$.
Let $I$ be an Iwahori subgroup fixed by a chamber in the apartment, and let $K_{0}$ be the special maximal compact subgroup corresponding to the special point $x_{0}$. The group $I$ has an Iwahori decomposition with respect to $B$ and $M$:
$I=(I\cap N^{-})(I\cap M)(I\cap N).$
Set $I_{\alpha}=I\cap w_{\alpha}Iw_{\alpha}$ for $\alpha\in\Delta$, and $I_{\widetilde{\alpha}-1}=I\cap w_{\widetilde{\alpha}-1}Iw_{\widetilde{\alpha}-1}$ for $\widetilde{\alpha}\in\widetilde{\Delta}$.
For each affine root $\psi$, let $U_{\psi}$ denote the corresponding root subgroup. Then $I$ is generated by $M_{0}$ and the subgroups $U_{\psi}$ with $\psi>0$. In particular,
$I\cap N=\prod_{\alpha\in\Phi^{+}}U_{\alpha},\qquad I\cap N^{-}=\prod_{\alpha\in\Phi^{+}}U_{-\alpha+1},$
and $I\cap M=M_{0}$.
For any $w\in\widetilde{W}$, denote by $q(w)$ the index $[IwI:I]$. For $\alpha\in\Phi_{\mathrm{aff}}$, let $q_{\alpha}=[U_{\alpha-1}:U_{\alpha}]$. Then one has
[…]
Let $\chi$ be an unramified character of $M$ (trivial on $M_{0}$), extended to a character of $B$ by making it trivial on $N$. Define the unramified principal series representation $I(\chi)=\operatorname{Ind}_{B}^{G}\chi$ (unnormalized induction). The space of Iwahori-fixed vectors $I(\chi)^{I}$ has a basis $\{\phi_{w,\chi}\mid w\in W\}$ described in [2, Prop. 2.1]. Concretely, for $p\in B$, $w^{\prime}\in W$, $b\in I$,
$\phi_{w,\chi}(pw^{\prime}b)=\begin{cases}\chi(p)&\text{if }w^{\prime}=w,\\ 0&\text{otherwise}.\end{cases}$
From the results of [7, Theorem 3.1] one deduces the action of the characteristic functions of the double cosets IsIIsI ($s\in S_{\mathrm{aff}}$) in the Iwahori–Hecke algebra $\mathcal{H}(G,I)$. The following corollary records the formulas we need, adapted to unnormalized induction.
Corollary 4.2.
For $w\in W$, $\alpha\in\Delta$, and $\widetilde{\alpha}\in\widetilde{\Delta}$, we have in $\mathcal{H}(G,I)$ (cf. [7][3.2], adapted to unnormalized induction):
$\operatorname{ch}_{Iw_{\alpha}I}*\phi_{w,\chi}$ $=\begin{cases}q_{\alpha+1}\phi_{ww_{\alpha},\chi}+(q_{\alpha+1}-1)\phi_{w,\chi}&\text{if }w\alpha\in\Phi^{-},\\[4.0pt] \phi_{ww_{\alpha},\chi}&\text{if }w\alpha\in\Phi^{+},\end{cases}$ (1)
$\operatorname{ch}_{Iw_{\widetilde{\alpha}-1}I}*\phi_{w,\chi}$ $=\begin{cases}\chi(a_{w\widetilde{\alpha}})\phi_{ww_{\widetilde{\alpha}},\chi}&\text{if }w\widetilde{\alpha}\in\Phi^{-},\\[4.0pt] q_{\widetilde{\alpha}}\chi(a_{w\widetilde{\alpha}})\phi_{ww_{\widetilde{\alpha}},\chi}+(q_{\widetilde{\alpha}}-1)\phi_{w,\chi}&\text{if }w\widetilde{\alpha}\in\Phi^{+}.\end{cases}$ (2)
5 Representation theory of the finite Lie group
In this section let us suppose that $G$ is a reductive Lie group defined over $\mathbb{F}_{q}$, $B$ is its Borel, $T$ is the maximal torus contained in $B$, and $U$ is its unipotent radical. $F$ is the geometric Frobenius endomorphism. The material of this section comes from Chapter 4 of the book of Geck and Jacon [4].
Definition 5.1.
The ring $H=\operatorname{Hom}_{\mathbb{C}[G^{F}]}(\operatorname{Ind}_{B^{F}}^{G^{F}}\mathbf{1},\operatorname{Ind}_{B^{F}}^{G^{F}}\mathbf{1})^{\mathrm{opp}}$ (In order to simplify the terminology, from now on, we will omit ”opp”. )is called the Iwahori–Hecke algebra of $G$.
[4] tells us the relations between irreducible subrepresentations of the principal Harish-Chandra series and simple $H$-modules.
Theorem 5.2.
The functor $\Theta:M\longmapsto\operatorname{Hom}_{G^{F}}(\operatorname{Ind}_{B^{F}}^{G^{F}}\mathbf{1},M)$ induces a bijection from irreducible subrepresentations of the principal Harish-Chandra series to the simple Hecke algebra modules up to isomorphism. The action of this $H$-module is $h*f=f\circ h$ for $h\in H$ and $f\in\operatorname{Hom}_{G^{F}}(\operatorname{Ind}_{B^{F}}^{G^{F}}\mathbf{1},M)$.
We can go one step further using Frobenius reciprocity:
$\operatorname{Hom}_{G^{F}}(\operatorname{Ind}_{B^{F}}^{G^{F}}\mathbf{1},M)=\operatorname{Hom}_{B^{F}}(\mathbf{1},M)=M^{B^{F}}.$
Hence if we have a finite dimensional representation $V$ of $G^{F}$, and we assume further that every irreducible component of $V$ is a subrepresentation of $\operatorname{Ind}_{B^{F}}^{G^{F}}\mathbf{1}$, then we can take $B^{F}$-invariant vectors of $V$ to make it into an $H$-module. Also note that ${Ind}_{B^{F}}^{G^{F}}\mathbf{1}=\mathbb{C}[G^{F}/B^{F}]$.
Then for an arbitrary element $n$ in the normalizer $N_{G^{F}}(T)$, we can define an element of $H$:
$T_{n}:\operatorname{Ind}_{B^{F}}^{G^{F}}\mathbf{1}\longrightarrow\operatorname{Ind}_{B^{F}}^{G^{F}}\mathbf{1}$
which sends $xB^{F}$ to the sum of all cosets $yB^{F}$ such that $x^{-1}y\in B^{F}nB^{F}$. We also define $q_{w}=[B^{F}wB^{F}:B^{F}]$ for an arbitrary element $w$ in the Weyl group. Then we have:
Theorem 5.3.
$\{T_{w}\mid w\in W\}$ forms a basis of the Iwahori–Hecke algebra $H$. If we further assume that $S$ is the set of simple reflections, then $\{T_{w}\mid w\in W\}$ forms a basis of $H$, and $H$ has the following presentation:
$H=\left\langle T_{s}\;(s\in S)\;\middle|\;\underbrace{T_{s}T_{t}T_{s}\cdots}_{m_{st}\text{ factors}}=\underbrace{T_{t}T_{s}T_{t}\cdots}_{m_{st}\text{ factors}},\;(T_{s}-q_{s})(T_{s}+1)=0\right\rangle.$
Finally, if we have a finite dimensional representation $V=\bigoplus_{i}V_{i}$ where each $V_{i}$ is irreducible, and we assume further that every irreducible component of $V$ is a subrepresentation of $\operatorname{Ind}_{B^{F}}^{G^{F}}\mathbf{1}$, then we can take the $B^{F}$-fixed vectors of $V$. We obtain an $H$-module $V^{B^{F}}=\bigoplus_{i}V_{i}^{B^{F}}$. It is well known that the Steinberg representation is a 1-dimensional representation on which $T_{\alpha}$ (where $\alpha$ is a simple root) acts as 1-1. Hence by Theorem 5.1, we obtain the following corollary:
Corollary 5.4.
The multiplicity of the Steinberg representation in $V$ equals $\dim_{\mathbb{C}}\{v\in V\mid T_{\alpha}v=-v\text{ for \ every \ simple \ root\ }\alpha\}$.
We should also note that by definition, $q_{s_{\alpha}}=q_{\alpha+1}$ for a simple root $\alpha$, where $q_{\alpha+1}$ is the element defined in Section 4.2.
6 Relations between finite Lie group and $p$-adic group
Two Hecke algebras occur in the previous sections, so we will compare these two algebras in this section. To simplify terminology, let $H=\operatorname{End}_{K/K_{+}}(\operatorname{Ind}_{I/K_{+}}^{K/K_{+}}(1))$, and let $H(K,I)$ be the finite Hecke algebra of the $p$-adic group (i.e., $I$-bi-invariant functions which are supported on $K$).
Lemma 6.1.
There is an isomorphism $\phi$ between $H$ and $H(K,I)$ satisfying $\phi(T_{s_{\alpha}})=\operatorname{ch}_{Is_{\alpha}I}$ for every simple root $\alpha$.
Proof.
This follows directly from the presentations of the two algebras. ∎
Lemma 6.2.
Suppose $G$ is an unramified reductive $p$-adic group with hyperspecial subgroup $K$ and Iwahori subgroup $I$. Let $(\pi,V)$ be an irreducible Iwahori-spherical representation of $G$. As a representation of the $p$-adic group, $V$ has an action coming from $H(K,I)$. As a representation of the finite Lie group $K/K_{+}$, according to Section 5, after taking $I/K_{+}$-fixed vectors, $V^{I}$ is also an $H$-module. Then the two actions coincide.
Proof.
Here we emphasize that the measure on $G$ is normalized such that the volume of $I$ is 1. By Lemma 3.5, every irreducible component of $V^{K_{+}}$ is a subrepresentation of $\operatorname{Ind}_{I/K_{+}}^{K/K_{+}}(1)$.
In order to show that these two actions coincide, we just need to check the generators $T_{s_{\alpha}}$ and $\operatorname{ch}_{Is_{\alpha}I}$. More concretely, we need to show the following equality: for every element $v\in V^{I}$
$T_{s_{\alpha}}v=\operatorname{ch}_{Is_{\alpha}I}v.$
Assume that the representatives of $Is_{\alpha}I/I$ are $\{u\mid u\in\Lambda\}$. Then By the discussion of the action of $T_{s_{\alpha}}$ in Section 5, and after applying Frobenius reciprocity, we have:
$\operatorname{Hom}_{K/K_{+}}(\operatorname{Ind}_{I/K_{+}}^{K/K_{+}}(1),V^{K_{+}})=\operatorname{Hom}_{I/K_{+}}(1,V^{K_{+}})=V^{I}$
For an element $\in V^{I}, vv$ corresponding to the function $f_{v}$, where $f_{v}(gI/K_{+})=gv$, hence
$T_{s_{\alpha}}v=f_{v}\circ T_{s_{\alpha}}(I/K_{+})$
$f_{v}\bigl(\sum_{u\in\Lambda}uI/K_{+}\bigr).$
$\sum_{u\in\Lambda}uv,$
$\text{RHS}=\int_{Is_{\alpha}I}xv\,dx=\sum_{u\in\Lambda}uv.$
Hence LHS = RHS. ∎
7 The classification of irreducible representations containing Steinberg
7.1 Main Theorem
Theorem 7.1.
Suppose that $G$ is an unramified reductive $p$-adic group, $K$ is its hyperspecial subgroup, and $I$ is its Iwahori subgroup. If an irreducible representation $(\pi,V)$ of $G$ contains Steinberg, then it must be an Iwahori-spherical representation. Moreover, for every unramified principal series of $G$, there is exactly one subquotient that contains Steinberg, and we will give a concrete construction in the proof of the theorem.
Proof.
The first statement follows from Corollary 3.4.
We now prove the second statement. Suppose that $(\pi,V)$ is an Iwahori-spherical irreducible representation, and it is a subquotient of the unramified principal series $I(\chi)$.
$V^{K_{+}}\subset I(\chi)^{K_{+}}$ is a sub-$K/K_{+}$ representation. Since $V$ is a subquotient of $I(\chi)$, the space $V^{K_{+}}$ is a subquotient of $I(\chi)^{K_{+}}$. The latter is a principal series representation of the finite reductive group $K/K_{+}$; hence all its irreducible constituents lie in the principal Harish-Chandra series. Consequently, every irreducible component of $V^{K_{+}}$ lies in the principal series.
By the representation theory of finite groups of Lie type [3], we can take the $I/K_{+}$ fixed points of $V^{K_{+}}$ (where $I/K_{+}$ is the Frobenius fixed point of a Borel subgroup of the finite Lie group $K/K_{+}$) to make it into a module of the Iwahori–Hecke algebra of the finite Lie group. So we need to study $V^{I}=(V^{K_{+}})^{I/K_{+}}$.
By Lemma 6.2, the two Hecke algebras have the same action; we can make $V^{I}$ an $H$-module of the finite Lie group.
In order to find the multiplicity of the Steinberg representation in $V^{K_{+}}$, we need to compute the dimension of the space
$W=\{v\in V^{I}\mid\operatorname{ch}_{Is_{\alpha}I}v=-v\text{ for \ every \ simple \ root\ }\alpha\}.$
We use the basis mentioned in Section 4; suppose that $v=\sum_{w\in W}c(w)\phi_{w,\chi}$. We prove the following lemma:
Lemma 7.2.
There is a unique (up to scalar) non-zero element of $I(\chi)$ that belongs to the space $W$.
Proof.
For a fixed simple root $\alpha$, we can pair the elements of $W$ as $\{(w,ws_{\alpha})\mid w\alpha\text{ is positive}\}$. According to Corollary 4.2:
$\operatorname{ch}_{Is_{\alpha}I}(c(w)\phi_{w,\chi}+c(ws_{\alpha})\phi_{ws_{\alpha},\chi})$
$(c(w)+c(ws_{\alpha})q_{\alpha+1}-c(ws_{\alpha}))\phi_{ws_{\alpha},\chi}+(c(ws_{\alpha})q_{\alpha+1})\phi_{w,\chi}.$
Since $v$ lies in $W$, we see that
$c(w)=c(ws_{\alpha})q_{\alpha+1}$
and
$c(ws_{\alpha})=c(w)+c(ws_{\alpha})q_{\alpha+1}-c(ws_{\alpha}).$
Hence we conclude that
$c(w)+q_{\alpha+1}c(ws_{\alpha})=0.$
Since $\alpha$ is arbitrary, we finish the uniqueness of the lemma via induction:
Let $c(1)=1$. From the above, we obtain $c(s_{\alpha})=-q^{-1}_{\alpha+1}$. Suppose we already know $c(w)$ for every $l(w)=m$. Now consider an element $u\in W$ with length $m+1$. There exists a simple root $\alpha$ such that $l(us_{\alpha})=l(u)-1$; then by the above argument $c(u)$ can be determined by $c(us_{\alpha})$, hence by the induction hypothesis we prove uniqueness. In other words, $\dim W\leq 1$. Next, let us prove existence: set $c(w)=(-1)^{l(w)}([IwI:I])^{-1}$. Note that the basic property of BN-pairs says: If for $w\in W$ and a simple reflection $s_{\alpha}$ we have $l(ws_{\alpha})=l(w)+1$, then $IwI:I][Is_{\alpha}I:I]=[Iws_{\alpha}I:I]$. This property tells us that there exists a non-zero element in $W$. Since if $w\in W$ such that $w(\alpha)>0$, for a simple root $\alpha$. then $l(ws_{\alpha})=l(w)+1$, using the above property we can see that $c(ws_{\alpha})q_{\alpha+1}+c(w)=0$. Hence $\sum_{w\in W}(-1)^{l(w)}[IwI:I]^{-1}\phi_{w,\chi}$ is a non-zero element in W. ∎
Returning to the proof of our main theorem, we see that the dimension of $W$ is 1, and we conclude that there is only one subquotient of the series $I(\chi)$ which contains Steinberg.
Finally, we give a construction of this subquotient. Denote the element in $W$ by $v_{0}$. Let $V$ be the subrepresentation of $I(\chi)$ generated by $v_{0}$, and let $V_{0}$ be the maximal subrepresentation of $V$ which does not contain $v_{0}$. Let’s us show that such $V_{0}$ exists: Since $I(\chi)$ is of finite length, suppose that the Jordan-Hölder sequence of $I(\chi)$ is
$0\subset U_{1}\subset U_{2}....\subset U_{m}=I(\chi)$
And by the previous argument, we can assume that the irreducible subquotient $U_{i}/U_{i-1}$ contains a Steinberg. Then $v_{0}\in U_{i}^{I}$, hence $V\subset U_{i}$, and since $v_{0}\notin U_{i-1}^{I}$, that means $V\cap U_{i-1}\neq V$, and there exists a natural G-injection $V/(V\cap U_{i-1})\xrightarrow{}U_{i}/U_{i-1}$. Since $U_{i}/U_{i-1}$ is irreducible and $V/(V\cap U_{i-1})\neq 0$, we can see that $V/(V\cap U_{i-1})$ is isomorphic to $U_{i}/U_{i-1}$. Hence we can take $V_{0}=V\cap U_{i-1}$, such $V_{0}$ exists. Then $V/V_{0}$ is the unique subquotient of $I(\chi)$ that contains the Steinberg representation. ∎
It is well known that two principal series $I(\chi)$ and $I(w\chi)$ have the same Jordan-Hölder factors (Where w is an element of the Weyl group). And if $\chi$ and $\chi^{\prime}$ are not in the same $W$ -orbit, then $I(\chi)$ and $I(\chi^{\prime})$ do not have a common Jordan-Hölder factor. Hence we get a corollary:
Corollary 7.3.
There is a bijection between $W$-orbits of unramified characters and irreducible representations containing the Steinberg in its hyperspecial subgroup.

### 31 Mar 2026 v2

@@
 Abstract
-This paper addresses Question 1 posed by Dipendra Prasad in his recent problem list: classify all irreducible smooth representations $\pi$ of an unramified reductive $p$-adic group $G$ (over a characteristic zero local field) such that the space of $K_{+}$-fixed vectors $V^{K_{+}}$ (viewed as a representation of the finite reductive group $K/K_{+}$) contains the Steinberg representation.
-We prove that any such representation $\pi$ must be Iwahori-spherical, hence a subquotient of some unramified principal series $I(\chi)=\operatorname{Ind}_{B}^{G}\chi$.
-By a detailed analysis of the action of the Iwahori–Hecke algebra on the Iwahori-fixed space $V^{I}$, we show that in each principal series $I(\chi)$ there exists exactly one irreducible subquotient containing Steinberg, and we give an explicit construction of this subquotient.
-Hence this can give us a bijection between W-orbits of unramified characters and irreducible representations containing Steinberg in its hyperspecial subgroup.
-This classification is richer than Prasad’s expectation: besides twists of the Steinberg representation and generic unramified representations, every principal series contributes a unique subquotient with the required property.
-The proof relies on establishing an isomorphism between the finite Iwahori–Hecke algebra of the finite Lie group $K/K_{+}$ and the finite Hecke algebra $H(K,I)$ of the $p$-adic group (compactly supported $I$-bi-invariant functions on $K$), and comparing their actions on $V^{I}$.
-Keywords: $p$-adic groups, unramified reductive groups, Steinberg representation, Iwahori subgroup, Hecke algebras, unramified principal series, Moy–Prasad filtration.
+This paper addresses Question 1 posed by Dipendra Prasad in his recent problem list [15]: classify all irreducible smooth representations $\pi$ of an unramified reductive $p$-adic group $G$ (over a characteristic zero local field) such that the space of $K_{+}$-fixed vectors $V^{K_{+}}$ (viewed as a representation of the finite reductive group $K/K_{+}$) contains the Steinberg representation.
+Prasad initially expected that such representations should be twists of the Steinberg representation or generic unramified representations.
+In a subsequent private communication, he refined the expectation: they should be precisely the unique generic irreducible subquotients of unramified principal series.
+We prove that any such $\pi$ must be Iwahori-spherical, hence a subquotient of some unramified principal series $I(\chi)=\operatorname{Ind}_{B}^{G}\chi$.
+By a detailed analysis of the Iwahori-Hecke algebra action on the Iwahori-fixed space $V^{I}$, we show that each principal series $I(\chi)$ contains exactly one irreducible subquotient whose $K_{+}$-fixed vectors contain the Steinberg representation, and we give an explicit construction of this subquotient.
+Consequently, we obtain a bijection between $W$-orbits of unramified characters and irreducible representations containing the Steinberg representation in their hyperspecial subgroup.
+The proof is self-contained and uses only elementary Hecke algebra methods.
+In particular, the reduction from the $K$-action on $K_{+}$-fixed vectors to modules over the finite Hecke algebra $H(K,I)$ is spelled out in detail, a step that has not been made explicit in the literature.
 1 Introduction
@@
 Classify all irreducible representations of $G(F)$ containing the Steinberg representation of $G(\mathbb{F}_{q})$ as a $K$-subrepresentation (i.e., such that $V^{K_{+}}$ contains the Steinberg representation of $K/K_{+}$ as a constituent).
-Prasad suggested that such representations should fall into two families: (1) twists of the Steinberg representation of $G(F)$; (2) generic, unramified representations (i.e., irreducible spherical principal series).
-However, our investigation reveals a more nuanced picture: every unramified principal series $I(\chi)=\operatorname{Ind}_{B}^{G}\chi$ (where $B$ is a Borel subgroup and $\chi$ an unramified character of a maximal torus) contains exactly one irreducible subquotient that contains the Steinberg representation.
-In particular, the class of representations satisfying Prasad’s condition is strictly larger than his expectation; each principal series contributes a unique such subquotient, hence this can give us a bijection between W-orbits of unramified characters and irreducible representations containing Steinberg in its hyperspecial subgroup.
-The main steps of our argument are as follows.
-Let $(\pi,V)$ be an irreducible smooth representation of $G(F)$ with $V^{K_{+}}$ containing the Steinberg representation.
-By the depth-zero theory of Moy–Prasad [8, 9], $\pi$ must be of depth zero.
-Using Lemma 3.2 and the Borel–Casselman theorem [1], we deduce that $\pi$ is Iwahori-spherical, i.e., $V^{I}\neq 0$ for an Iwahori subgroup $I\subset K$.
-Hence $\pi$ is a subquotient of some unramified principal series $I(\chi)$.
-The space $V^{K_{+}}$ is a representation of the finite reductive group $K/K_{+}$; by Lemma 3.1 all its irreducible constituents lie in the principal Harish-Chandra series, i.e., are subrepresentations of $\operatorname{Ind}_{I/K_{+}}^{K/K_{+}}\mathbf{1}$.
-Taking $I/K_{+}$-fixed points yields $V^{I}$, which carries two natural actions: one from the Hecke algebra $H_{\mathrm{fin}}=\operatorname{End}_{K/K_{+}}(\operatorname{Ind}_{I/K_{+}}^{K/K_{+}}\mathbf{1})$ (coming from finite group theory), and one from the finite Hecke algebra $H(K,I)$ of the $p$-adic group (compactly supported $I$-bi-invariant functions on $K$).
-In Section 6 we prove that these two actions coincide on $V^{I}$ (Lemma 6.2), allowing us to transfer known results about the Steinberg representation from finite groups to the $p$-adic setting.
-The multiplicity of the Steinberg representation in $V^{K_{+}}$ equals the dimension of the space
-$W=\{v\in V^{I}\mid T_{\alpha}v=-v\text{ for\ every \ simple \ root \ }\alpha\},$
-where $T_{\alpha}$ are the standard generators of the Hecke algebra.
-Using Casselman’s basis $\{\phi_{w,\chi}\}$ of $I(\chi)^{I}$ and Lansky’s formulas for the action of affine reflections [7] (adapted to unnormalized induction), we solve this eigenvalue problem in $I(\chi)^{I}$.
-Lemma 7.2 shows that the solution space is one‑dimensional; a non‑zero vector is explicitly given by
+Prasad initially suggested that such representations should fall into two families: (1) twists of the Steinberg representation of $G(F)$; (2) generic unramified representations (i.e., irreducible spherical principal series).
+In a subsequent private communication, he refined this expectation: the representations in question should be precisely the unique generic irreducible subquotients of unramified principal series.
+The present paper confirms this refined expectation and gives a complete classification. Our main results are as follows.
+  • Existence and uniqueness.
+For every unramified principal series $I(\chi)=\operatorname{Ind}_{B}^{G}\chi$ (where $B$ is a Borel subgroup and $\chi$ an unramified character of a maximal torus), there exists exactly one irreducible subquotient whose $K_{+}$-fixed vectors contain the Steinberg representation.
+We give an explicit construction of this subquotient via the Iwahori-Hecke algebra.
+  • Parametrization.
+The construction yields a bijection between $W$-orbits of unramified characters (where $W$ is the relative Weyl group) and irreducible representations of $G(F)$ whose $K_{+}$-fixed vectors contain the Steinberg representation.
+Our proof is self-contained and uses only elementary methods from the theory of Iwahori-Hecke algebras. The main steps are:
+  1. 1.
+    Show that any representation satisfying Prasad’s condition must be Iwahori-spherical (hence a subquotient of some $I(\chi)$) using the Moy–Prasad theory of depth-zero representations and the Borel–Casselman theorem.
+  2. 2.
+    Reduce the problem to an eigenvalue problem on the Iwahori-fixed space $V^{I}$ by comparing two actions on $V^{I}$: one coming from the finite Hecke algebra $H(K,I)$ of the $p$-adic group, and the other from the Iwahori–Hecke algebra of the finite Lie group $K/K_{+}$.
+We prove that these two actions coincide (Lemma 6.2), a step that has not been spelled out clearly in the literature.
+  3. 3.
+    Solve the eigenvalue problem using Casselman’s basis $\{\phi_{w,\chi}\}$ of $I(\chi)^{I}$ and Lansky’s formulas for the action of affine reflections.
+The unique (up to scalar) solution is given explicitly by
     $v_{0}=\sum_{w\in W}(-1)^{\ell(w)}[IwI:I]^{-1}\phi_{w,\chi}.$
-Consequently $\dim W=1$, so the Steinberg representation occurs with multiplicity one in $V^{K_{+}}$.
-Moreover, letting $V$ be the subrepresentation of $I(\chi)$ generated by $v_{0}$ and $V_{0}$ the maximal subrepresentation of $V$ not containing $v_{0}$, the quotient $V/V_{0}$ is the unique irreducible subquotient of $I(\chi)$ that contains Steinberg.
-This completes the classification.
-The paper is organized as follows.
-Section 2 recalls the Moy–Prasad filtration and unrefined minimal $K$-types.
-Section 3 analyses the cuspidal support of the irreducible components of $V^{K_{+}}$ and proves that a representation containing Steinberg must be Iwahori-spherical.
-Section 4 collects facts about unramified principal series: Casselman’s basis and Lansky’s action of affine reflections.
+  4. 4.
+    Show that the subrepresentation generated by $v_{0}$ contains a unique irreducible subquotient with the desired property.
+And we also show that this subquotient is the unique generic subquotient in a unramified principal series.
+Relation to previous work
+After the first version of this paper appeared on arXiv, we were informed by Manish Mishra and Maarten Solleveld that related results exist in the literature.
+We are very grateful to them for their careful reading and for pointing out the connections.
+  • Manish Mishra [9] has studied generic representations in $L$-packets.
+In Theorem 12 of [9], he proved the existence and uniqueness theorem under the framework of unramified L-packets.
+In the split case, Theorem 3 of his joint work with Basudev Pattanayak [10] gives an explicit Hecke-algebra model for the unique generic constituent of an unramified principal series.
+These results, which are part of a more general framework for $L$ packages, imply our classification in a broader context.
+  • Maarten Solleveld [16] developed a comprehensive theory of principal series representations for quasi-split reductive $p$-adic groups.
+In Theorem 3.4 of [16], he established that the condition used by Prasad is equivalent to the representation being generic.(To see that one has to know as well the reduction steps from the K-action on K+-invariant vector to modules over H(K,I), which is contained in this article.
+) His work, which is much more general, provides a conceptual explanation for our results.
+Professor Solleveld also kindly pointed out that the reduction steps from the $K$-action on $K_{+}$-fixed vectors to modules over H(K,I)H(K,I)—which we work out in detail—are not explicitly written elsewhere, and that this part of our work can be regarded as an original contribution.
+We are deeply indebted to them for their generosity and for helping us locate our work within the larger framework.
+Original contributions of this paper
+While the existence and uniqueness of generic subquotients of unramified principal series are known from general theory, our work provides the following new elements:
+  1. 1.
+    Explicit construction.
+We give an explicit formula for a basis vector $v_{0}$ in the Iwahori-fixed space that generates the desired subquotient.
+This explicit description is new and may be useful for further computations .
+  2. 2.
+    Self-contained elementary proof. Our proof uses only Hecke algebra calculations and does not rely on the more advanced machinery of $L$-packets or the local Langlands correspondence.
+  3. 3.
+    Detailed reduction steps.
+The comparison between the finite Hecke algebra $H(K,I)$ and the Iwahori–Hecke algebra of the finite Lie group $K/K_{+}$ (Lemma 6.2) is made completely explicit.
+As noted by Solleveld, this step, has not been written out clearly in the literature.
+Organization of the paper
+Section 2 recalls the Moy–Prasad filtration and the theory of unrefined minimal $K$-types.
+Section 3 analyzes the cuspidal support of the irreducible components of $V^{K_{+}}$ and proves that any representation containing the Steinberg representation must be Iwahori-spherical.
+Section 4 collects necessary facts about unramified principal series, including Casselman’s basis and Lansky’s formulas for the action of affine reflections.
 Section 5 reviews the representation theory of finite Lie groups, focusing on the principal series and the associated Iwahori–Hecke algebra.
-Section 6 establishes the isomorphism between the finite Hecke algebra and the $p$-adic finite Hecke algebra $H(K,I)$, and proves that their actions on $V^{I}$ coincide.
-Section 7 contains the core lemma and the proof of the main theorem.
+Section 6 establishes an isomorphism between the finite Hecke algebra of the finite group and the finite Hecke algebra $H(K,I)$ of the $p$-adic group, and proves that their actions on $V^{I}$ coincide.
+Section 7 contains the core lemma and the proof of the main theorem, including the explicit construction of the subquotient.
+Finally, we include a corollary giving a bijection between $W$-orbits of unramified characters and the classified representations.
+Section 8 provides readers with another view of this result.
 2 Unrefined minimal $K$-types and the Moy–Prasad filtration
@@
 For a point $x$ in the Bruhat–Tits building of $G/k$, let $\mathcal{P}_{x}=\mathcal{P}_{x}\cap G(k)$ be the corresponding parahoric subgroup of $G(k)$.
-Moy and Prasad associate to $x$ a decreasing filtration $\{\mathcal{P}_{x,r}\}_{r\geq 0}$ of $\mathcal{P}_{x}$ by open compact subgroups, defined as follows: for a non-negative real number $r$, let $\mathcal{P}_{x,r}$ be the subgroup generated by certain congruence subgroups of the maximal torus and by the root subgroups $U_{\psi}$ for all affine roots $\psi$ satisfying $\psi(x)\geq r$; see [8, 2.6] for the precise definition.
+Moy and Prasad associate to $x$ a decreasing filtration $\{\mathcal{P}_{x,r}\}_{r\geq 0}$ of $\mathcal{P}_{x}$ by open compact subgroups, defined as follows: for a non-negative real number $r$, let $\mathcal{P}_{x,r}$ be the subgroup generated by certain congruence subgroups of the maximal torus and by the root subgroups $U_{\psi}$ for all affine roots $\psi$ satisfying $\psi(x)\geq r$; see [11, 2.6] for the precise definition.
 One sets $\mathcal{P}_{x,0}=\mathcal{P}_{x}$ and $\mathcal{P}_{x,r+}=\bigcup_{s>r}\mathcal{P}_{x,s}$.
@@
 There is an associated filtration of the Lie algebra $\mathfrak{g}$.
-Let $\mathfrak{g}_{x,r}$ be the $\mathfrak{o}$-lattice in $\mathfrak{g}$ obtained from the smooth $\mathfrak{o}$-group scheme corresponding to $\mathcal{P}_{x}$; its definition is given in [9, 3.2].
+Let $\mathfrak{g}_{x,r}$ be the $\mathfrak{o}$-lattice in $\mathfrak{g}$ obtained from the smooth $\mathfrak{o}$-group scheme corresponding to $\mathcal{P}_{x}$; its definition is given in [12, 3.2].
 Similarly, one defines a filtration $\{\mathfrak{g}_{x,r}^{*}\}_{r\in\mathbb{R}}$ of the dual $\mathfrak{g}^{*}$ by
 $\mathfrak{g}_{x,r}^{*}=\{X\in\mathfrak{g}^{*}\mid X(Y)\in\mathfrak{o}\text{ for all }Y\in\mathfrak{g}_{x,-r+}\},$
-see [8, 3.5]. The pairing
+see [11, 3.5]. The pairing
 $\mathfrak{g}_{x,-r}^{*}/\mathfrak{g}_{x,(-r)+}^{*}\times\mathfrak{g}_{x,r}/\mathfrak{g}_{x,r+}\to\mathfrak{f}$
-is non-degenerate and $\mathbf{M}_{x}(\mathfrak{f})$-equivariant, where $\mathbf{M}_{x}$ is the reductive quotient of the special fibre of the integral model attached to $x$; cf. [9, 3.7].
+is non-degenerate and $\mathbf{M}_{x}(\mathfrak{f})$-equivariant, where $\mathbf{M}_{x}$ is the reductive quotient of the special fibre of the integral model attached to $x$; cf. [12, 3.7].
 For a character $\chi$ of $\mathcal{P}_{x,r}$ trivial on $\mathcal{P}_{x,r+}$ (with $r>0$), one can realize $\chi$ as a character of the abelian group $\mathcal{P}_{x,r}/\mathcal{P}_{x,r+}$.
-Via the isomorphism $\mathcal{P}_{x,r}/\mathcal{P}_{x,r+}\cong\mathfrak{g}_{x,r}/\mathfrak{g}_{x,r+}$ of [9, 3.8] and the above duality, such a character corresponds to a coset
+Via the isomorphism $\mathcal{P}_{x,r}/\mathcal{P}_{x,r+}\cong\mathfrak{g}_{x,r}/\mathfrak{g}_{x,r+}$ of [12, 3.8] and the above duality, such a character corresponds to a coset
 $X+\mathfrak{g}_{x,-r+}^{*}\in\mathfrak{g}_{x,-r}^{*}/\mathfrak{g}_{x,-r+}^{*}.$
@@
 Proof.
-For a fixed Iwahori subgroup $I$ of $G$ , By [6, Theorem 8.4.10(2)], we know that $I/I_{+}$ is the minimal Levi of $K/K_{+}$.
+For a fixed Iwahori subgroup $I$ of $G$ , By [7, Theorem 8.4.10(2)], we know that $I/I_{+}$ is the minimal Levi of $K/K_{+}$.
 Since every finite Lie group is quasi-split, $I/I_{+}$ is a torus, and $I/K_{+}$ is the Borel containing this torus.
@@
 $\phi_{w,\chi}(pw^{\prime}b)=\begin{cases}\chi(p)&\text{if }w^{\prime}=w,\\ 0&\text{otherwise}.\end{cases}$
-From the results of [7, Theorem 3.1] one deduces the action of the characteristic functions of the double cosets IsIIsI ($s\in S_{\mathrm{aff}}$) in the Iwahori–Hecke algebra $\mathcal{H}(G,I)$.
+From the results of [8, Theorem 3.1] one deduces the action of the characteristic functions of the double cosets IsIIsI ($s\in S_{\mathrm{aff}}$) in the Iwahori–Hecke algebra $\mathcal{H}(G,I)$.
 The following corollary records the formulas we need, adapted to unnormalized induction.
@@
 There is a bijection between $W$-orbits of unramified characters and irreducible representations containing the Steinberg in its hyperspecial subgroup.
+Remark: In Theorem 3.4 of [16], Professor Solleveld established that the condition used by Prasad is equivalent to the representation being generic.(To see that one has to know as well the reduction steps from the K-action on K+-invariant vector to modules over H(K,I), which is contained in this article.
+)
+8 Another Perspective: generic representation
+Theorem 8.1.
+The representation constructed above is the unique generic subquotient in the unramified principal series $I(\chi)$.
+Proof.
+It is well known that there is exactly one subquotient of $I(\chi)$ is generic.
+And the work of Moeglin and Waldspurger [13] told us an irreducible representation $(\pi,V)$ is generic if and only if the Wavefront set of $\pi$ is the regular orbit.
+From Example 2.3.8 of Okada’s paper [14] we can see that if $V^{K_{+}}$ contains Steinberg, then the geometry wavefront set of this representation must be the regular nilpotent orbit.
+Hence the representation that we construct above is the unique generic subquotient.
+∎